OpenStudy (anonymous):

|sin(x)| = 1/2 solve the equation on the interval [0,2pi)

OpenStudy (anonymous):

possibilities are \[\sin(x)=\frac{1}{2}\] or \[\sin(x)=-\frac{1}{2}\] do you know how to find \(x\) ?

OpenStudy (anonymous):

so I need it in radians, I'd like to know how to do that :)

OpenStudy (anonymous):

find the point on the unit circle where the second coordinate is \(\frac{1}{2}\) on the last page of the attached cheat sheet. there will be two such places. then look at the angle repeat the process for \(-\frac{1}{2}\)

OpenStudy (anonymous):

as for "radians" the trig functions are functions of numbers. as functions of angles, they only correspond to the functions of numbers if the angles are measured in radians, so forget degrees unless you are solving a triangle and you need to measure in degrees

OpenStudy (anonymous):

so my answer is going to be : {pi/6, 7pi/6 }

OpenStudy (anonymous):

if you open the cheat sheet, you will see the point \((\frac{\sqrt3}{2},\frac{1}{2})\) and right next to it you will see \(\frac{\pi}{6}\) so that is one answer

OpenStudy (anonymous):

yep, that's what I did.

OpenStudy (anonymous):

you should have a total of 4 solutions, two for \[\sin(x)=\frac{1}{2}\]and two more for \[\sin(x)=-\frac{1}{2}\] you got one for each, but there is one more for each

OpenStudy (anonymous):

look directly to the left of the first one you got \(x=\frac{\pi}{6}\) and you will see another ordered pair \((-\frac{\sqrt3}{2},\frac{1}{2})\) beside that is the number \(\frac{5\pi}{6}\) so that is also a solution

OpenStudy (anonymous):

oh yeah, right \[\pi/6, 5\pi/6, 7\pi/6, 11\pi/6 \]

OpenStudy (anonymous):

thank you, I got it now :)

OpenStudy (anonymous):

yw